### v1

@@ Abstract
Early-stage 3D brain tumor segmentation from magnetic resonance imaging (MRI) scans is crucial for prompt and effective treatment. However, this process faces the challenge of precise delineation due to the tumors’ complex heterogeneity. Moreover, energy sustainability targets and resource limitations, especially in developing countries, require efficient and accessible medical imaging solutions. The proposed architecture, a Lightweight 3D ATtention U-Net with Parallel convolutions, LATUP-Net, addresses these issues. It is specifically designed to reduce computational requirements significantly while maintaining high segmentation performance. By incorporating parallel convolutions, it enhances feature representation by capturing multi-scale information. It further integrates an attention mechanism to refine segmentation through selective feature recalibration. LATUP-Net achieves promising segmentation performance: the average Dice scores for the whole tumor, tumor core, and enhancing tumor on the BraTS 2020 dataset are , , and , and on the BraTS 2021 dataset, they are , , and , respectively. Hausdorff distance metrics further indicate its improved ability to delineate tumor boundaries. With its significantly reduced computational demand using only M parameters, about times fewer than other state-of-the-art models, and running on a single V100 GPU, LATUP-Net stands out as a promising solution for real-world clinical applications, particularly in settings with limited resources. Investigations into the model’s interpretability, utilizing gradient-weighted class activation mapping and confusion matrices, reveal that while attention mechanisms enhance the segmentation of small regions, their impact is nuanced. Achieving the most accurate tumor delineation requires carefully balancing local and global features. The code is available at https://qyber.black/ca/code-bca.
@@
Brain tumors, particularly gliomas, represent some of the most lethal forms of cancer due to their inherent complexity, and high degree of variability among patients. Gliomas, classified into high-grade and low-grade gliomas, encompass various tumoral structures, including the enhancing tumor, necrotic tumor core, and edema [1]. Magnetic resonance imaging (MRI) is the standard imaging technique for diagnosing gliomas. MRI’s central role in glioma imaging underscores the importance of precise segmentation for comprehensive tumor treatment and sparing healthy tissue, thereby influencing diagnosis, treatment planning, and prognostication. The subjectivity and labor-intensiveness of manual segmentation underscore the necessity for more efficient and accurate automatic segmentation techniques [2].
In response to the limitations of manual tumor segmentation, advances have been made towards fully automatic methods, specifically employing deep learning. Convolutional neural networks (CNNs) have demonstrated significant achievements in several computer vision domains, including, but not limited to, image segmentation, classification, and object detection. They are commonly utilized in medical imaging due to their characteristics of gaining local receptive fields, allowing weight sharing, and reducing parameters by pooling [3, 4]. However, despite the laudable progress in deep-learning-based strategies, brain tumor segmentation remains challenging due to tumor heterogeneity, anatomical variations, inconsistencies in imaging protocols, diverse and irregular tumor appearances, imaging artifacts, and inherent uncertainties in medical images. In addition, current deep architectures, particularly those employing 3D convolutions, have complex structures, high computing costs, and risk of overfitting, particularly with limited data [5]. These issues are especially significant in real-world applications like MRI image segmentation [3].
The increasing reliance on medical images and clinician-to-patient ratio challenges in low- and middle-income countries necessitate high precision, intelligent automated systems for timely and accurate medical diagnoses [6]. Moreover, energy sustainability targets and resource limitations, especially in geographic regions with limited computational resources, favor solutions requiring less computational resources. Hence, advances in medical image segmentation are geared towards exploring and developing models that are not only efficient but also resource-conscious, making them suitable for a wide range of computing environments [7].
Deep neural networks (DNNs) extract features that contain both, essential and redundant, information, ensuring a comprehensive understanding of input images. For example, inception blocks [8] utilize parallel convolutional paths with varying kernel sizes and a max-pooling operation to capture features at different scales and abstraction levels starting from the same input. However, each path in an inception block operates independently from the start, without a shared feature extraction stage. This may lead to redundancy and an increase in the number of parameters, as each path may learn similar features. While this redundancy may enrich the feature set for DNN training, there is a trend towards reducing it towards more resource-efficient approaches.
Our proposed architecture, the LATUP-Net (Lightweight 3D ATtention U-Net with Parallel convolutions), is a novel U-Net variant with fewer parameters and comparable results to state-of-the-art architectures. It balances network depth and width, which is crucial for achieving optimal segmentation performance in lightweight models. Our work introduces an innovative parallel convolutions (PC) approach, derived from the inception block [8], within the first encoder block to tackle the challenge of feature redundancy, and efficiently harness diverse features at varying scales and orientations. After an initial shared embedded convolutional layer, the model processes the data through parallel paths with distinct kernel sizes followed by immediate pooling, capturing multi-scale spatial features. Compared to the inception block [8], this reduces feature redundancy and parameter count by leveraging a shared feature extraction stage. It simplifies the parallel block’s paths and reduces convolutional depth before concatenation, but maintains diverse features while keeping the parameter count lower than that of inception blocks. This strategy addresses computational and memory constraints common in 3D image processing tasks and meets the demand for more resource-efficient models in real-world applications, where limitations often exist.
This paper further bridges the gap between the need for efficient models and the allure of attention mechanisms. We propose a lightweight model that significantly reduces computational requirements while maintaining high segmentation performance. We also investigate the intricate dynamics of attention mechanisms, which leads us to conclude that the attention mechanism effectively highlights tumor features but may overlook broader contextual information essential for accurate segmentation. LATUP-Net’s performance is evaluated using the per-sample Dice similarity coefficient (DSC) and the Hausdorff distance (HD95). Results show comparable performance to state-of-the-art models while reducing the number of model parameters from over million in the state-of-the-art nnU-Net [9] to million. This offers benefits in terms of sustainability and accessibility.
The contributions of our work are:
We propose LATUP-Net, an efficient, novel 3D U-Net variant, which incorporates parallel convolutions and attention mechanisms to achieve high brain tumor segmentation performance at highly reduced computational costs. Empirically, we show that LATUP-Net achieves significantly better segmentation accuracy than U-Net with half as many parameters and considerably fewer parameters than other state-of-the-art models with comparable performance.
We demonstrate the efficacy of parallel convolutions in capturing features at different scales and orientations, leading to a richer representation of the input data and enhanced segmentation results. The parallel convolutions significantly improve efficiency and segmentation accuracy by incorporating a shared embedded convolution and a subsequent max-pooling operation following each parallel path.
We engage in a detailed exploration of attention mechanisms. The crux of our investigation lies in assessing whether integrating attention invariably enhances segmentation or if there are scenarios where local context captured by convolutions suffices. We integrate an attention mechanism in the model and interpret its behavior using gradient-weighted class activation mapping (Grad-CAM) [10] and confusion matrix analysis. This reveals a nuanced trade-off: while attention effectively emphasizes salient features within tumor regions, it leads to an oversight of the broader contextual information essential for accurate segmentation. A more balanced approach that considers local details and global contextual cues could improve the model’s segmentation performance, giving us a fuller picture of the tumor’s structure.
@@ 2 Related Work
To analyze existing methods for brain tumor segmentation, we consider three perspectives: convolutional neural network (CNN) capacity, lightweight models, and attention mechanisms
2.1 CNN Models Capacity
@@
While 3D convolutions make better use of spatial information, they also require more computational power and memory. To address this issue, Chen et al. [12] developed a memory-efficient solution, while preserving most of the volumetric information, by introducing a decoupled 3D U-Net model. It relies on separating a 3D convolution into sequential 2D and 1D convolutions and creating three parallel branches of these separated convolutions, one for each orthogonal view (axial, sagittal, and coronal) for the 1D convolution direction. The suggested model achieved competitive results while demonstrating high efficiency when tested on the BraTS 2018 dataset. While local and global features are necessary for making decisions during segmentation, low-level feature gradients (such as those containing information about boundaries, edges, lines, or dots) converge to zero as one proceeds deeper into the model. To address this, Wang et al. [13] proposed a TransBTS architecture that effectively embeds a transformer into a 3D U-Net model. To begin with, local feature maps are extracted using a 3D CNN encoder. The extracted features are transmitted through a transformer to capture global features. Afterward, the decoder incorporates the local and global features during the upsampling process to produce the segmentation result. Though this model requires more computational resources, it has shown promising results on the BraTS 2019 and 2020 datasets, achieving competitive or higher Dice score performance compared with state-of-the-art 3D models.
Traditional U-Net architectures perform exceptionally well on semantic segmentation tasks [14]. Nevertheless, such structures lack strategies to extract global feature information [15, 16]. To address this, the inception module [8] and a densely connected module [17] were added to the U-Net architecture by Zhang et al. [18]. Each inception module in the network uses , , and convolution kernels to acquire multi-scale information. Their method performs admirably in segmenting images of lung tissue, blood arteries, and brain tumors. However, while this approach improves the representation power of the model, it also increases the number of parameters, which heightens the risk of overfitting and limits its effectiveness in scenarios with limited training data. Thus, given the constraints of low resource funding in hospitals, it becomes crucial to strike a balance between processing efficiency and network size through the deployment of lightweight networks.
@@ 2.3 Attention Mechanism
The attention mechanisms explored include Squeeze-and-Excitation (SE) [22], the Convolutional Block Attention Module (CBAM) [24], Efficient Channel Attention (ECA) [25], and Residual Squeeze-and-Excitation (RSE) [26]. We further introduce a modified variant of SE where the fully connected (dense) layers are replaced by a 3D convolutional layer. We also experiment with combined mechanisms such as fusing CBAM and SE. The motivation behind combining CBAM and SE is to leverage the strengths of both. CBAM’s ability to focus on pertinent spatial and channel features and SE’s capacity to recalibrate channel-wise features may enhance the model’s ability to capture complex interdependencies in the data. Another approach is designed to exploit the strengths of convolutions while adhering to the principles of the SE mechanism. Replacing the dense layer in SE with a 3D convolution layer (SE-3D) is aimed at maintaining the spatial information of the input tensor and capturing local spatial correlations, while simultaneously maintaining the ability to recalibrate channel-wise features.
@@ 3 The LATUP-Net Architecture
Here, we explain the components of our LATUP-Net architecture, illustrated in Figure 1, a lightweight variant of the original U-Net [11] with fewer parameters intended for the semantic segmentation of 3D brain tumors. Moreover, LATUP-Net utilizes multi-scale parallel convolutions (see Section 3.1) and channel attention on multi-modal data fusion (see Section 3.2).

### v2

@@ Abstract
Early-stage 3D brain tumor segmentation from magnetic resonance imaging (MRI) scans is crucial for prompt and effective treatment. However, this process faces the challenge of precise delineation due to the tumors’ complex heterogeneity. Moreover, energy sustainability targets and resource limitations, especially in developing countries, require efficient and accessible medical imaging solutions. The proposed architecture, a Lightweight 3D ATtention U-Net with Parallel convolutions, LATUP-Net, addresses these issues. It is specifically designed to reduce computational requirements significantly while maintaining high segmentation performance. By incorporating parallel convolutions, it enhances feature representation by capturing multi-scale information. It further integrates an attention mechanism to refine segmentation through selective feature recalibration. LATUP-Net achieves promising segmentation performance: the average Dice scores for the whole tumor, tumor core, and enhancing tumor on the BraTS 2020 dataset are , , and , and on the BraTS 2021 dataset, they are , , and , respectively. Hausdorff distance metrics further indicate its improved ability to delineate tumor boundaries. With its significantly reduced computational demand using only M parameters, about times fewer than other state-of-the-art models, and running on a single NVIDIA GeForce RTX3060 12 GB GPU, LATUP-Net requires just GFLOPs. This makes it a promising solution for real-world clinical applications, particularly in settings with limited resources. Investigations into the model’s interpretability, utilizing gradient-weighted class activation mapping and confusion matrices, reveal that while attention mechanisms enhance the segmentation of small regions, their impact is nuanced. Achieving the most accurate tumor delineation requires carefully balancing local and global features. The code is available at https://qyber.black/ca/code-bca.
@@
Brain tumors, particularly gliomas, are among the most lethal forms of cancer due to their inherent complexity and high variability among patients. Gliomas, classified into high-grade and low-grade, consist of different tumor regions, including the enhancing tumor, necrotic core, and surrounding edema [1]. Magnetic resonance imaging (MRI) is the standard method for diagnosing gliomas, and precise segmentation of these tumors is crucial for effective treatment planning. Accurate segmentation helps clinicians differentiate tumor tissue from healthy tissue, which directly influences diagnosis, treatment strategies, and prognosis. However, manual delineation of tumor regions across MRI slices is both time-consuming and labor-intensive, with the accuracy highly dependent on the clinician’s expertise and subjective thresholding. This subjectivity, combined with the effort involved, underscores the need for efficient and accurate automatic segmentation techniques for brain tumors [2].
In this work we introduce a novel lightweight deep learning architecture, LATUP-Net (Lightweight 3D ATtention U-Net with Parallel Convolutions). It significantly reduces the computational resources needed while maintaining state-of-the-art brain tumor segmentation performance, as demonstrated on BraTS 2020 [1, 3, 4] and BraTS 2021 [1]. We incorporate parallel convolutions in the first encoder block of a U-Net architecture, inspired by the inception block [5]. This reduces feature redundancy and parameter count by sharing an initial feature extraction stage across multiple convolutional paths, followed by pooling operations to capture multi-scale spatial features. The design harnesses diverse features efficiently while maintaining a lower parameter count than traditional inception blocks. We further add an extension mechanism, which generally yields smaller performance improvements, but also does not add substantial computational costs.
Deep learning, particularly convolutional neural networks (CNNs), has revolutionized medical image analysis by providing powerful tools for segmentation, classification, and object detection. CNNs are widely used in medical imaging due to their ability to automatically extract hierarchical features, making them well-suited for complex tasks like brain tumor segmentation [6, 7]. Among these architectures, U-Net has become a leading model for medical image segmentation and has demonstrated remarkable performance across various tasks. The U-Net architecture has been widely adopted, receiving over citations as of 2023, reflecting its profound impact on medical image segmentation research [8, 9]. To meet the increasing demands of clinical applications, U-Net-based models have evolved by enhancing their network structures, incorporating new modules, and expanding to 3D networks [10, 11, 12]. However, improving segmentation accuracy often comes at the cost of increased model complexity and parameters [13]. Existing networks require large numbers of parameters and computational resources, which poses challenges for real-world use in resource-limited environments [14]. To address this challenge, lightweight networks have been developed to reduce model parameters and computational resource requirements without compromising performance [6, 15]. These models employ strategies such as full convolution and deep separable convolution to achieve high segmentation accuracy while minimizing resource usage, making them more suitable for deployment in resource-constrained environments. In our specific approach we exploit parallel convolutions to recuce computational resources requierd for training and inference.
In particular, in MRI scans with small tumor lesions, attention mechanisms have been employed to improve results to focus on a specific region and its features. Such integrating can improve the accuracy of small tumor segmentation [16]. Common lightweight attention methods for medical image semantic segmentation include BAM [17], CBAM [18], and Squeeze-and-Excitation (SE) [19]. BAM uses atrous convolution to achieve a larger receptive field, CBAM combines spatial and channel attention separately, and SE focuses on channel attention to address feature loss during convolutional pooling. However, using both spatial and channel attention increases computational complexity. The Squeeze-and-Excitation mechanism is chosen for its efficient feature extraction through channel attention while maintaining computational efficiency, making it suitable for small-scale segmentation tasks [19].
We explore the integration of attention mechanisms within LATUP-Net to improve segmentation of tumor regions. While attention mechanisms effectively emphasize tumor-specific features, our investigation highlights a trade-off: attention may focus too narrowly on these features and overlook broader contextual information, which is also essential for accurate segmentation. By balancing local detail and global context, LATUP-Net achieves performance comparable to state-of-the-art models while drastically reducing computational demands.
In summary, the contributions of our work are:
We introduce LATUP-Net, an efficient 3D U-Net variant that combines parallel convolutions and attention mechanisms to achieve high segmentation accuracy at a fraction of the computational cost of existing models.
We demonstrate the effectiveness of parallel convolutions in capturing multi-scale features, resulting in a richer, more efficient representation. They optimize feature diversity and computational efficiency by leveraging a shared initial convolution followed by distinct paths and pooling operations.
We critically examine the role of attention mechanisms in segmentation, using Grad-CAM [20] and confusion matrix analysis to assess their impact. Our findings show that while attention enhances focus on tumor features, a balanced approach considering both local detail and global context improves overall segmentation performance.
@@ 2 Related Work
To analyze existing methods for brain tumor segmentation, we consider three perspectives: convolutional neural networks (CNN) and their variants, lightweight models, and attention mechanisms
2.1 CNN and Related Models
@@
While 3D convolutions make better use of spatial information, they also require more computational power and memory. To address this issue, Chen et al. [22] developed a memory-efficient solution, while preserving most of the volumetric information, by introducing a decoupled 3D U-Net model. It relies on separating a 3D convolution into sequential 2D and 1D convolutions and creating three parallel branches of these separated convolutions, one for each orthogonal view (axial, sagittal, and coronal) for the 1D convolution direction. The suggested model achieved competitive results while demonstrating high efficiency when tested on the BraTS 2018 dataset. While local and global features are necessary for making decisions during segmentation, low-level feature gradients (such as those containing information about boundaries, edges, lines, or dots) converge to zero as one proceeds deeper into the model. To address this, Wang et al. [23] proposed a TransBTS architecture that effectively embeds a transformer into a 3D U-Net model. To begin with, local feature maps are extracted using a 3D CNN encoder. The extracted features are transmitted through a transformer to capture global features. Afterward, the decoder incorporates the local and global features during the upsampling process to produce the segmentation result. Zhu et al. [24] propose a BTS method that combines deep semantic features and edge features for semantic feature extraction and fusion from multimodal MRI. The method uses the Swin Transformer for semantic feature extraction, shifted patch tokenization for training efficiency, and an Edge Spatial Attention Block (ESAB) for feature enhancement. Though both models require more computational resources, it has shown promising results on BraTS 2018 to BraTS 2020, achieving competitive or higher Dice score performance compared with state-of-the-art 3D models.
Traditional U-Net architectures perform exceptionally well on semantic segmentation tasks [8]. Nevertheless, such structures lack strategies to extract global feature information [25, 26]. To address this, the inception module [5] and a densely connected module [27] were added to the U-Net architecture by Zhang et al. [28]. Each inception module in the network uses , , and convolution kernels to acquire multi-scale information. Their method performs admirably in segmenting images of lung tissue, blood arteries, and brain tumors. Meanwhile, the transformers’ self-attention mechanism automatically brings in global information but lacks the inductive bias, so it does not obtain sufficient fine-grained features. Thereby, combining transformers and CNNs may leverage the strengths of both. However, this combination often neglects lesion boundaries, which are essential for accurate segmentation. To address this issue, Xu et al. [11] integrated the Swin-T network with a dual-path feature inference module to enhance the original Swin-T network, resulting in improved edge segmentation performance for cranial tumors. Zhu et al. [10] used skip connections to integrate multi-level edge fusion features, derived from the sparse dynamic encoder, into the decoder, therefore improving the transmission of spatial edge information and further refining the network’s segmentation performance.
While these approaches improve the representational power of the models, they also increase the number of parameters, heightening the risk of overfitting and limiting effectiveness in scenarios with limited training data. Therefore, given the constraints of low resource funding in many hospitals, it is crucial to strike a balance between processing efficiency and network size through the development of lightweight networks.
@@ 2.2 Lightweight Models
In a recent study, Zhu et al. [32] proposed a CNN-based model for brain tumor segmentation. Their approach combines three modules that use multimodal, spatial, and boundary information. This method examined the overall spatial aspects of the image allowing it to accurately acquire the tumor’s location and how it relates to other tissues in MRI scans. The proposed model proved to be more effective and efficient than other existing state of the arts methods.
@@ 3 The LATUP-Net Architecture
Here, we explain the components of our LATUP-Net architecture, illustrated in Figure 1 and Table 1, a lightweight variant of the original U-Net [21] with fewer parameters intended for the semantic segmentation of 3D brain tumors. Moreover, LATUP-Net utilizes multi-scale parallel convolutions (see Section 3.1) and channel attention on multi-modal data fusion (see Section 3.2).
@@
The LATUP-Net architecture is intentionally designed to minimize the dependence on complex ensembling or additional computational resources. In alignment with best practices outlined in [9], we focus on delivering an efficient, lightweight model that demonstrates its innovations without depending on confounding performance boosters. This approach ensures fair comparisons and highlights the true impact of our architectural choices. By combining proven techniques, such as multi-scale parallel convolutions and lightweight attention mechanisms, LATUP-Net balances the need for high segmentation performance with reduced computational complexity, making it suitable for resource-constrained environments [35].
@@
Layer | Input Shape | Layer Type | Stride | Output Shape | Parameters |
Input Layer | (128, 128, 128, 3) | Input | - | (128, 128, 128, 3) | 0 |
Conv3D_1 (enc1_pc_embed) | (128, 128, 128, 3) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 2624 |
Conv3D_2 (enc1_pc_1_conv) | (128, 128, 128, 32) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 1056 |
Conv3D_3 (enc1_pc_2_conv) | (128, 128, 128, 32) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 27680 |
Conv3D_4 (enc1_pc_3_conv) | (128, 128, 128, 32) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 128032 |
MaxPooling3D_1 (enc1_pc_1_maxpool) | (128, 128, 128, 32) | MaxPooling3D | (2,2,2) | (64, 64, 64, 32) | 0 |
MaxPooling3D_2 (enc1_pc_2_maxpool) | (128, 128, 128, 32) | MaxPooling3D | (2,2,2) | (64, 64, 64, 32) | 0 |
MaxPooling3D_3 (enc1_pc_3_maxpool) | (128, 128, 128, 32) | MaxPooling3D | (2,2,2) | (64, 64, 64, 32) | 0 |
Concatenate (enc1_pc_concat) | (64, 64, 64, 32) | Concatenate | - | (64, 64, 64, 96) | 0 |
SE Layer_1 (enc2_SE_mult) | (64, 64, 64, 96) | Squeeze and Excitation | - | (64, 64, 64, 96) | 2304 |
Conv3D_5 (enc2_conv1) | (64, 64, 64, 96) | Conv3D (64 filters) | (1,1,1) | (64, 64, 64, 64) | 165952 |
InstanceNorm_1 (enc2_instance_norm) | (64, 64, 64, 64) | Instance Normalization | - | (64, 64, 64, 64) | 128 |
Conv3D_6 (enc2_conv2) | (64, 64, 64, 64) | Conv3D (64 filters) | (1,1,1) | (64, 64, 64, 64) | 110656 |
Dropout_1 (enc2_dropout) | (64, 64, 64, 64) | Dropout | - | (64, 64, 64, 64) | 0 |
MaxPooling3D_4 (enc2_maxpool) | (64, 64, 64, 64) | MaxPooling3D | (2,2,2) | (32, 32, 32, 64) | 0 |
SE Layer_2 (enc3_SE_mult) | (32, 32, 32, 64) | Squeeze and Excitation | - | (32, 32, 32, 64) | 1024 |
Conv3D_7 (enc3_conv1) | (32, 32, 32, 64) | Conv3D (128 filters) | (1,1,1) | (32, 32, 32, 128) | 221312 |
InstanceNorm_2 (enc3_instance_norm) | (32, 32, 32, 128) | Instance Normalization | - | (32, 32, 32, 128) | 256 |
Conv3D_8 (enc3_conv2) | (32, 32, 32, 128) | Conv3D (128 filters) | (1,1,1) | (32, 32, 32, 128) | 442496 |
Dropout_2 (enc3_dropout) | (32, 32, 32, 128) | Dropout | - | (32, 32, 32, 128) | 0 |
MaxPooling3D_5 (enc3_maxpool) | (32, 32, 32, 128) | MaxPooling3D | (2,2,2) | (16, 16, 16, 128) | 0 |
SE Layer_3 (bn_SE_mult) | (16, 16, 16, 128) | Squeeze and Excitation | - | (16, 16, 16, 128) | 4096 |
UpSampling3D_1 (dec3_upsample) | (16, 16, 16, 128) | UpSampling3D | (2,2,2) | (32, 32, 32, 128) | 0 |
Conv3D_9 (dec3_conv1) | (32, 32, 32, 128) | Conv3D (128 filters) | (1,1,1) | (32, 32, 32, 128) | 131200 |
InstanceNorm_3 (dec3_instance_norm) | (32, 32, 32, 128) | Instance Normalization | - | (32, 32, 32, 128) | 256 |
Concatenate (dec3_concat) | (32, 32, 32, 128) | Concatenate | - | (32, 32, 32, 256) | 0 |
Conv3D_10 (dec3_conv2) | (32, 32, 32, 256) | Conv3D (128 filters) | (1,1,1) | (32, 32, 32, 128) | 884864 |
UpSampling3D_2 (dec2_upsample) | (32, 32, 32, 128) | UpSampling3D | (2,2,2) | (64, 64, 64, 128) | 0 |
Conv3D_11 (dec2_conv1) | (64, 64, 64, 128) | Conv3D (64 filters) | (1,1,1) | (64, 64, 64, 64) | 65600 |
InstanceNorm_4 (dec2_instance_norm) | (64, 64, 64, 64) | Instance Normalization | - | (64, 64, 64, 64) | 128 |
Concatenate (dec2_concat) | (64, 64, 64, 64) | Concatenate | - | (64, 64, 64, 128) | 0 |
Conv3D_12 (dec2_conv2) | (64, 64, 64, 128) | Conv3D (64 filters) | (1,1,1) | (64, 64, 64, 64) | 221248 |
UpSampling3D_3 (dec1_upsample) | (64, 64, 64, 64) | UpSampling3D | (2,2,2) | (128, 128, 128, 64) | 0 |
Conv3D_13 (dec1_conv1) | (128, 128, 128, 64) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 16416 |
Concatenate (dec1_concat) | (128, 128, 128, 32) | Concatenate | - | (128, 128, 128, 64) | 0 |
Conv3D_14 (dec1_conv2) | (128, 128, 128, 64) | Conv3D (32 filters) | (1,1,1) | (128, 128, 128, 32) | 55328 |
Conv3D_15 (prob_map) | (128, 128, 128, 32) | Conv3D (4 filters) | (1,1,1) | (128, 128, 128, 4) | 132 |
Total Parameters | 3,069,060 |